OpenStudy (brittany16):

What number needs to be eliminated to solve this equation? -8/X=72

OpenStudy (anonymous):

8x9 so, 8

OpenStudy (brittany16):

no go away

OpenStudy (lgbasallote):

first cross multiply which gives you -8 = 72x now divide both sides by 72 which gives you x = -1/9

OpenStudy (anonymous):

-(8)/x=(8)9 cancel 8. -1/x=9 multiply both sides by x -1(x)/x=9x x cancel divide by nine x=-1/9. By using this way you eliminate a number before multiplying like good mathematicians do.